### v1

@@ Abstract
In recent years, climate extremes such as floods have created significant environmental and economic hazards for Australia, causing damage to the environment and economy and losses of human and animal lives. An efficient method of forecasting floods is crucial to limit this damage. Techniques for flood prediction are currently based on hydrological, and hydrodynamic (physically-based) numerical models. Machine learning methods that include deep learning offer certain advantages over conventional physically based approaches, including flexibility and accuracy. Deep learning methods have been promising for predicting small to medium-sized climate extreme events over a short time horizon; however, large flooding events present a critical challenge. We present an ensemble-based machine learning approach that addresses large-scale extreme flooding challenges using a switching mechanism motivated by extreme-value theory for long-short-term-memory (LSTM) deep learning models. We use a multivariate and multi-step time-series prediction approach to predict streamflow for multiple days ahead in the major catchments of Australia. The ensemble framework also employs static information to enrich the time-series information, allowing for regional modelling across catchments. Our results demonstrate enhanced prediction of streamflow extremes, with notable efficacy for large flooding scenarios in the selected Australian catchments. Through comparative analysis, our methodology underscores the potential for deep learning models to revolutionise flood forecasting across diverse regions.
@@ 1 Introduction
In Australia, flooding is among the most devastating and costly natural disasters in terms of economic loss [1, 2], damage to natural habitats and environment, and loss of human and animal lives [3, 4]. Additionally, Fernandez et al. [5] highlighted the psychological impact of floods and suggested that floods can cause post-traumatic stress disorder, depression and anxiety, and long-term mental health implications for flood victims. It is the rare and extreme floods which have the most far-reaching consequences for both humans and the environment. Between 1998 and 2022 multiple extreme floods occurred in NSW and Queensland causing between \$900 million and \$3.4 billion losses. As well as causing fatalities, economic impacts continued for years afterwards.
@@
Machine learning methods have been increasingly used to tackle time series prediction problems for a wide range of domains [6, 7, 8, 9] and are extensively used in hydrology [10, 11]. Machine learning methods can forecast streamflow at various lead-times, and show promise for hydrological forecasts of streamflow and floods [12, 13, 14, 15, 16, 17]. Hence, the literature motivates the use of machine learning models for extreme climate events such as floods in Australia.
@@
A major challenge in hydrology is the ’regional modelling problem’, which is about the usage of either a model or a set of models to provide spatially continuous hydrological simulations across large areas (e.g., regional, continental, global). Regional models must learn and encode the similarities and differences of various catchment characteristics (such as topology, geometry, climate, ecology, etc.) and use them to meaningfully predict hydrological behaviour. Often this is done using ancillary data (e.g. soil maps, remote sensing, digital elevation maps, etc.) to map the similarity between catchments [18]. Kratzert et al. [19] showed how an long short-term memory (LSTM) [20, 21] model could be used to develop a single (regional) hydrologic model for hundreds of sites, concatenating the static data to the time-series data as inputs. LSTM models have outperformed traditional hydrological models for streamflow prediction in ungauged basins [22]. The compression of the static information used in hydrologic modeling can improve the integration of the static information with the time series data [23]. Ensemble learning methods use multiple machine learning models and combine their predictions for improved performance. In this case, the ensemble models complement the LSTM’s temporal information with the static features. Leontjeva et al. [24] showed that an LSTM model can extract information from temporal sequences, which can then be combined with ensemble methods like Extreme Value Machines and Random Forests to utilize static data. Miebs et al. [25] demonstrated that an ensemble with altered architectures of recurrent neural networks (RNNs) concatenated with dense networks could learn the temporal data with static supplementary information. Therefore, an ensemble machine learning model approach is promising, giving the opportunity for a regional approach to outperform a single-site approach calibrated per basin, which often suffers from inherent data shortage.
@@
Forecasting time series with extremes is very difficult as the extreme values are often treated as outliers to improve the generalisation of the model. Most research on extreme values focuses on using extreme value theory (EVT) [26] as applications in anomaly-detection (detecting that an extreme event is present), rather than being able to predict the magnitude of the extreme event. An example is the work by Galib et al. [27] who used neural networks to predict the parameters of an extreme value distribution which consequently could predict the highest value within a specific time window of the time-series data.
@@
Quantile regression models the conditional median (or other quantiles) of the response variable, as compared to linear regression, which estimates the conditional mean. Similarly, a quantile regression model training procedure minimizes the mean absolute error to estimate conditional quantiles of the response variable [28]. The quantile regression model is more robust to outliers in the response measurements and hence applicable to forecasting extremes [29, 30]. Quantile regression has been combined with machine learning models for several applications. Taylor et al. [31] introduced the quantile regression neural network, which offered a practical and flexible implementation compared to past work that involved linear or simple parametric non-linear models. Wang et al. [32] combined quantile regression with an LSTM model for load forecasting, and Pasche et al. [33] combined extreme value theory with quantile regression for a novel model which was applied to flood risk forecasting. Quantile regression found multiple applications in hydrology, such as in downscaling precipitation [34] and climate characteristics [35]. In relation to flood forecasting, Weerts et al. [36] applied quantile regression to assess the predictive uncertainty of rainfall-runoff and hydraulic forecasts. Cai and Reeve [30] addressed extreme and non-extreme value prediction via parametric quantile function models, showcasing differences among statistical models commonly used in extreme value prediction.
@@
The domain of extreme flood prediction, particularly through the integration of deep learning with quantile regression, presents a unique opportunity to enhance the accuracy of flood forecasts. Despite significant advancements in forecasting streamflow and rainfall-runoff modelling, there is a critical gap in relation to predicting extreme flood events. This gap could be addressed by combining deep learning with quantile regression, which so far has not been used widely. This will specifically allow for assessing predictive uncertainty in hydrological forecasts. This study introduces an ensemble-based deep learning framework, which uses a switching mechanism inspired by extreme value theory for deep learning models, employing a multivariate and multi-step time-series prediction approach to forecast streamflow for multiple days ahead in major Australian catchments. We compare the performance of various deep learning models, including LSTM, bi-directional-LSTM, and CNNs, across a large number of basins. Our framework also features an ensemble model that integrates quantile regression with deep learning, enriching the time series with static information to support regional modelling. We provide open-source code along with data so that the methodology can be extended to other catchments in other regions.
@@
The rest of the paper is organised as follows. In Section 2, we provide an overview of related topics, and in Section 3 we present our methodology. In Section 4, we provide the results, followed by a discussion in Section 5, and a conclusion in Section 6.
@@ 2.1 Hydrology
Surface water flooding is a result of accumulated rainfall covering a land surface since it cannot drain to natural or man-made drainage systems or watercourses due to them being at full capacity and overflowing [37]. In short duration or intense rainfall, flash flooding can arise from rapidly developing convective rainfall systems, which can be particularly dangerous in urban areas [38]. The prediction of flash floods for small to medium catchments remains a great challenge, as flash floods aren’t always caused by known meteorological phenomena [39]. Some factors that cause flash floods may include how quickly the storm is moving, the porosity of the soil, the steepness of the terrain, etc. Climate change is a big driver of extreme weather events that cause floods [40], and in particular rainfall events have been increasing in intensity [41].
@@
The most common hydrological forecasting models are combined hydrological/hydrodynamic and meteorological models [42]. Hydrological models mostly simulate rainfall-runoff relationships, while hydraulic models deal with channelflow and flow over surfaces. Meteorological rainfall forecast models are needed, as ultimately rainfall drives floods, but these need to quantify uncertainty [43]. The problem is that there are numerous sources of uncertainty, which limits forecasting ability [44]. The general approach to manage uncertainty has been to use ensemble forecasts. For example, Emerton et al. [45] used ensemble streamflow forecasting and featured a number of forecast scenarios to explicitly characterize forecast uncertainties. At a much larger scale, Alfieri et al. [46] presented a global flood awareness system (GloFAS) for detecting the probability of streamflow given warning thresholds by using ensemble models. However, Cloke et al. [47] outlined that ensemble models can be limited by computational speed and complexity. Ultimately, hydrological models are constrained by the amount of information available, such as the input of meteorological variables (e.g. flow, water level, rainfall) and the need to parameterise geological, topographic and soil variables [48].
As mentioned, most hydrological models are calibrated for a single catchment and the challenge is to extrapolate to different catchment across a region. In regionalisation approaches, model-dependent (parametric) and model-independent (non-parametric) approaches are the two primary strategies [18].
@@
The model-dependent approaches are those that assume a pre-defined hydrological model across all catchments, wherein the model parameters are to be estimated from the data. For example, different approaches can be used for regionalisation, either focusing on spatial proximity of or on hydrological similarity [49], and spatial proximity gave the best performance across catchments in France. However all approaches were outperformed by models calibrated to single catchments [49]. In a recent uncertainty focused approach, Prieto et al. [50] aimed at identifying not only model parameters, but different model structures in a regionalisation approach.
@@
Conversely, model-independent methods, non-parametric in nature, require no prior assumptions and are driven entirely by the data [22]. It identifies the mapping between ancillary data, catchment attributes and other inputs to predict streamflow. The most common of these are machine learning models such as neural networks, being most effective with large amounts of data, and allowing knowledge to be transferred between sites [51]. Several papers [52, 19, 53] have shown that LSTM approaches in particular can outperform a numerical conceptual hydrological models,even if they are calibrated individually for each catchment in a gauged setting [52].
2.2 Deep learning and ensemble models
The success of recursive neural networks (RNN) has been mostly demonstrated in language models [54], but it has also been prominent for modelling temporal sequences [55] and multistep ahead prediction [56]. Some of the major categories of time series forecasting (prediction) include one-step, multi-step and multi-variate time series prediction. The one-step ahead time series prediction refers to models that have a prediction horizon (step) of a single time unit, whereas multi-step predictions refers to models that have a prediction horizon of multiple time units. The challenge is in developing models that maintain good prediction accuracy as the prediction time step increases, since the chaotic nature and noise of time-series data makes this difficult [57, 56].
RNN models cannot learn long-term dependency, as these do not have "memory" to capture autocorrelation present in long streams of the data [58]. This limitation was addressed by the long short-term memory (LSTM) network [20] through the inclusion of internal memory cells that control the flow of information in the model. This has improved natural language processing (NLP) [59] and language translation [60] which involve longer-term dependencies, since they have data with long sequences, i.e a paragraph in a text or the data can be the entire chapter which needs to be translated.
@@ 2.2 Deep learning and ensemble models
Bi-directional RNN and LSTM (Bi-RNN, Bi-LSTM) further address the shortcomings of earlier counterparts, where only direct previous context states are used for determining future states. The bi-directional architecture consists of placing two independent RNN/LSTM layers together to allow information to be processed in both directions of the network, at every time step [61]. Accordingly, information is transferred from past to future and from future to past[62].
Convolutional neural networks (CNN) were designed to learn spatial hierarchies of features through multiple building blocks of convolution, pooling and fully connected layers. Convolution layer(s) in a CNN were originally designed for image classification purposes [63], but can also be used for temporal sequences, including time series prediction [64, 65, 56]. For example CNN has been used for precipitation downscaling [66].
Suppose $Y$ is the real-valued random value for the response, then $F_{Y}(y)=P(Y\leq y)$ represents the cumulative distribution function. The $\tau^{t}h$ quantile for $\tau\in(0,1)$ may be given by:
$q_{Y}(t)=F_{Y}^{-1}(\tau)=inf\{y:F_{Y}(y)\geq\tau\}$ (1)
Rather than using the squared error-based function, we define the tilted absolute value function as $\rho_{\tau}(m)$ for a particular quantile $\tau$.
$q_{Y}(\tau)=arg_{u}minE(\rho_{\tau}(Y-u))$ (2)
@@
As highlighted earlier in the introduction, deep learning and particular LTSM models can capture memory and therefore autocorrelation, as well as linear and non-linear behaviour in hydrological systems. LSTM models can learn long-term dependencies for temporal sequences and exhibit better performance for flood forecasting [67] when compared to simple neural networks. Although the CNN[64, 65] was designed for computer vision tasks, they also have shown remarkable results in time-series prediction. In particular, Chandra et al. [56] showed that CNN is competitive with LSTM for multi-step ahead time series prediction. Furthermore, Hussain et al. [68] implemented a one-dimensional (1D) CNN for one-step-ahead streamflow forecasting for three-time horizons (daily, weekly and monthly) in Gilgit River of Pakistan. Khosravi et al. [69] used CNN with meta-heuristic optimisation and reported superior predictive performance in daily stream-flow forecasting compared to related models.
@@
Similarly highlighted in the introduction, ensemble learning models combine multiple machine learning methods as base models to produce an optimal predictive model [70, 71]. Some of the prominent ensemble learning models are bagging and boosting via implementations such as random forests and adaboost [72]. In hydrology, ensemble models have also been applied extensively [11], where the main question has focused on how to construct the ensemble. Should this by averaging methods [73], bootstrapping [74] or stacking [75, 76]. Initially, Cannon et al. [77] investigated how climate change impacted streamflow using ensemble neural networks and multiple linear regression methods. Further work by Diks et al. [73] compared several ensemble methods such as equal weights averaging, and indicated that the Granger-Ramanathan averaging method most improved forecast accuracy for environmental systems. In contrast, Sharma et al. [74] demonstrated that a hybrid wavelet bootstrapped neural network provided better predictions than conventional machine learning models in forecasting daily river discharge. Tyralis et al. [75] took a different approach, to forecast short-term daily streamflow, by stacking ten different machine learning models. Furthermore, Li et al. [76] showed an example, which integrated stacked models such as support vector machines, elastic-net regression and extreme gradient boosting (XGBoost) model to forecast midterm stream-flow. However, bagging has been the predominant ensemble model until 2010 in hydrology, but more recently, boosting and stacking has dominated due to improved performance [11].

### v2

@@ Abstract
In recent years, climate extremes such as floods have created significant environmental and economic hazards for Australia. Deep learning methods have been promising for predicting extreme climate events; however, large flooding events present a critical challenge due to factors such as model calibration and missing data. We present an ensemble quantile-based deep learning framework that addresses large-scale streamflow forecasts using quantile regression for uncertainty projections in prediction. We evaluate selected univariate and multivariate deep learning models and catchment strategies. We implement a multistep time-series prediction model using the CAMELS dataset for selected catchments across Australia. The ensemble model employs a set of quantile deep learning models for streamflow determined by historical streamflow data. We utilise the streamflow prediction and obtain flood probability using flood frequency analysis and compare it with historical flooding events for selected catchments. Our results demonstrate notable efficacy and uncertainties in streamflow forecasts with varied catchment properties. Our flood probability estimates show good accuracy in capturing the historical floods from the selected catchments. This underscores the potential for our deep learning framework to revolutionise flood forecasting across diverse regions and be implemented as an early warning system.
@@ 1 Introduction
Flooding is among the most devastating and costly natural disasters in Australia in terms of economic loss [1, 2], damage to natural habitats and environment, and loss of human and animal lives [3, 4]. Fernandez et al. [5] highlighted the psychological impact of floods and suggested that floods can cause post-traumatic stress disorder, depression and anxiety, and long-term mental health implications for flood victims. It is the rare and extreme floods that have the most far-reaching consequences for both humans and the environment. Between 1998 and 2022, multiple extreme floods occurred in New South Wales (NSW) and Queensland, causing between \$900 million and \$3.4 billion losses. As well as causing fatalities, economic impacts continued for years afterwards.
@@
Machine learning methods have been increasingly used to tackle time series prediction problems for a wide range of domains [6, 7, 8, 9] and are extensively used in hydrology [10, 11]. Machine learning methods can forecast streamflow at various lead times, and show promise for hydrological forecasts of streamflow and floods [12, 13, 14, 15, 16, 17]. Hence, the literature motivates the use of machine learning models for extreme climate events such as floods in Australia.
@@
A major challenge in hydrology is the ’regional modelling problem’, which is about the usage of either a single model or a set of models to provide spatially continuous hydrological simulations across large areas (e.g., regional, continental, global). Regional models must learn and encode the similarities and differences of various catchment characteristics (such as topology, geometry, climate, ecology, etc.) and use them to meaningfully predict hydrological behaviour. Often this is done using ancillary data (e.g. soil maps, remote sensing, digital elevation maps, etc.) to map the similarity between catchments [18]. Kratzert et al. [19] showed how a Long Short-Term Memory (LSTM) [20, 21] deep learning model could be used to develop a single (regional) hydrologic model for hundreds of sites, concatenating the static data to the time-series data as inputs. LSTM models have outperformed traditional hydrological models for streamflow prediction in ungauged basins [22]. The compression of the static information used in hydrologic modelling can improve the integration of the static information with the time series data [23] using ensemble learning.
Ensemble learning models use multiple machine learning models and combine their predictions for improved performance [24, 25]. Leontjeva et al. [26] showed that an LSTM model can extract information from temporal sequences, which can then be combined with ensemble methods like Extreme Value Machines and Random Forests to utilise static data. Miebs et al. [27] demonstrated that an ensemble of Recurrent Neural Networks (RNNs) and Dense Networks could learn the temporal data with static supplementary information. Therefore, an ensemble learning models can enable a regional approach to potentially outperform a single-site approach calibrated per basin, which often suffers from inherent data shortage.
@@
Forecasting time series with extremes is very difficult, as the extreme values are often treated as outliers to improve the generalisation of the model. Most research on extreme values focuses on using Extreme Value Theory (EVT) [28] for detecting that an extreme event is present, rather than being able to predict the magnitude of the extreme event. Galib et al. [29] used simple neural networks to predict the parameters of an extreme value distribution, which consequently could predict the highest value within a specific time window of the time-series data.
@@
Quantile Regression models the conditional median (or other quantiles) of the response variable, whereas Linear Regression estimates the conditional mean. Similar to linear regression, a quantile regression model training procedure minimises the mean absolute error to estimate conditional quantiles of the response variable [30]. The quantile regression model is more robust to outliers in the response measurements and hence applicable to forecasting extremes [31, 32]. Quantile regression has been combined with machine learning models for several applications. Taylor et al. [33] introduced the quantile regression neural network, which offered a practical and flexible implementation compared to past work that involved linear or simple parametric non-linear models. Wang et al. [34] combined quantile regression with an LSTM model for load forecasting, and Pasche et al. [35] combined extreme value theory with quantile regression for a novel model which was applied to flood risk forecasting. Quantile regression found multiple applications in hydrology, such as in down-scaling precipitation [36] and climate characteristics [37]. In the case of flood forecasting, Weerts et al. [38] applied quantile regression to assess the predictive uncertainty of rainfall-runoff and hydraulic forecasts. Cai and Reeve [32] addressed extreme and non-extreme value prediction via parametric quantile function models, showcasing differences among statistical models commonly used in extreme value prediction.
@@
The domain of extreme flood prediction, particularly through the integration of deep learning with quantile regression, presents a unique opportunity to enhance the accuracy of flood forecasts. Despite significant advancements in forecasting streamflow and rainfall-runoff modelling, there is a critical gap concerning the prediction of extreme flood events that can enable the prediction of floods. This gap could be addressed by combining deep learning with quantile regression, which so far has not been used widely. This will specifically allow for assessing predictive uncertainty in hydrological forecasts. Furthermore, integration of quantile deep learning with ensemble learning [39, 40] can further address the problem of model uncertainty  quantification and enable uncertainty projection in forecasts.
In this study, we present an ensemble quantile deep learning framework that addresses large-scale streamflow and flood probability forecasts with uncertainty quantification in predictions. Our framework employs a multivariate and multi-step time-series prediction approach to forecast streamflow for multiple days ahead using the catchment attributes and meteorological (CAMELS) dataset [41] for selected catchments across Australian states. We evaluate selected univariate and multivariate deep learning models and catchment strategies, including individual and ensemble learning. We also employ static information to enrich the time-series information, allowing regional modelling across catchments. The ensemble model employs a set of quantile deep learning models for streamflow determined by historical streamflow data. We utilise the streamflow prediction and obtain flood probability using flood frequency analysis and compare it with historical flooding events for selected catchments. We compare the performance of selected deep learning models for selected basins across the Australian states. A major contribution of our study is the development of a quantile deep learning model within an ensemble model, which has major potential for hydrology. We provide open-source code along with data so that the methodology can be extended to other catchments in other regions.
@@
The rest of the paper is organised as follows. In Section 2, we provide an overview of related topics, and in Section 3 we present our methodology. Section 4 provides the results, followed by a discussion in Section 5, and a conclusion in Section 6.
@@ 2.1 Hydrological models
Surface water flooding is a result of accumulated rainfall covering a land surface since it cannot drain to natural or man-made drainage systems or watercourses due to them being at full capacity and overflowing [42]. In short-duration intense rainfall, flash flooding can arise from rapidly developing convective rainfall systems, which can be particularly dangerous in urban areas [43]. The prediction of flash floods for small to medium catchments remains a great challenge, as flash floods aren’t always caused by known meteorological phenomena [44]. Some factors that cause flash floods may include how quickly the storm is moving, the porosity of the soil, the steepness of the terrain, etc. Climate change is a big driver of extreme weather events that cause floods [45], and in particular rainfall events have been increasing in intensity [46].
@@
The most common hydrological forecasting models are combined hydrological/hydrodynamic and meteorological models [47]. Hydrological models mostly simulate rainfall-runoff relationships, while hydraulic models deal with channel flow over surfaces. Meteorological rainfall forecast models are needed, as ultimately rainfall drives floods, but these need to quantify uncertainty [48]. Several major sources of uncertainty limit forecasting ability [49] and the general approach to managing uncertainty has been to use ensemble forecasts. For example, Emerton et al. [50] used ensemble streamflow forecasting and featured several forecast scenarios to explicitly forecast uncertainties. At a much larger scale, Alfieri et al. [51] presented a Global Flood Awareness System (GloFAS) for detecting the probability of streamflow given warning thresholds by using ensemble models. However, Cloke et al. [52] outlined that ensemble models can be limited by computational speed and complexity. Ultimately, hydrological models are constrained by the amount of information available, such as the input of meteorological variables (e.g. flow, water level, rainfall) and the need to parameterise geological, topographic and soil variables [53].
@@
As mentioned, most hydrological models are calibrated for a single catchment, and the challenge is to extrapolate to different catchments across a region. In regionalisation approaches, model-dependent (parametric) and model-independent (non-parametric) approaches are the two primary strategies [18]. The model-dependent approaches assume a pre-defined hydrological model across all catchments, wherein the model parameters are estimated from the data. Different approaches can be used for regionalisation, focusing on either spatial proximity or hydrological similarity. In a study by Oudin et al. [54], spatial proximity gave the best performance across the selected catchments in France; however, the models calibrated to single catchments outperformed all the other approaches. In a recent uncertainty-focused approach, Prieto et al. [55] estimated not only the model parameters, but different model structures in a regionalisation approach.
@@
Conversely, model-independent methods, non-parametric in nature, require no prior assumptions and are driven entirely by the data [22]. It identifies the mapping between ancillary data, catchment attributes and other inputs to predict streamflow. The most common of these are machine learning models (such as neural networks) that have been most effective with large amounts of data and enable model knowledge to be transferred between sites [56]. Several studies [57, 19, 58] have shown that LSTM models, in particular, can outperform conceptual hydrological models, even if they are calibrated individually for each catchment in a gauged setting [57].
2.2 Deep learning models
The success of RNNs has been prominent in language models [59] and modelling temporal sequences [60], including multistep-ahead prediction [61]. The major categories of time series forecasting (prediction) include one-step, multi-step and multivariate time series prediction. The one-step-ahead prediction refers to models that have a prediction horizon (step) of a single time unit, whereas multi-step predictions refer to models that have a prediction horizon of multiple time units. Developing models that maintain good prediction accuracy as the prediction horizon increases is challenging [62, 61]. In certain cases, such as extreme value forecasting in climate-related problems, and cryptocurrency forecasting, the chaotic nature of the data and the missing features make a significant difference in model performance. RNN models had the problem of learning long-term dependencies in temporal data, due to the limited "memory" in capturing the autocorrelation present in long-term data streams [63]. The LSTM network [20] addressed this limitation through the inclusion of internal memory cells that control the flow of information in the model. This has improved Natural Language Processing (NLP) [64] and Machine Translation [65] which involve longer-term dependencies in text data, i.e a paragraph in a text can be the entire chapter that needs to be translated. LTSM models can capture memory and therefore autocorrelation, as well as linear and non-linear behaviour in hydrological systems. LSTM models can learn long-term dependencies for temporal sequences and exhibit better performance for flood forecasting [66] when compared to simple neural networks.
@@ 2.2 Deep learning models
Bidirectional RNNs [67] further addressed the shortcomings of earlier counterparts, where only direct previous context states are used for determining future states. The bidirectional architecture consists of placing two independent RNN layers together to allow information to be processed in both directions, at every time step [67]. Accordingly, information is transferred from past to future and from future to past time-steps, and bidirectional LSTM network implementation enhances conventional LSTM network [68]. The encoder-decoder LSTM [65] is an RNN originally designed to process sequence-to-sequence problems in text data processing, where the number of items in output and output sequences varies. In recent studies, the encoder-decoder LSTM has shown to be one of the best models, overtaking LSTM, CNN, and BD-LSTM for multi-step time series prediction problems [61, 69].
@@
CNNs [70, 71] have been designed to learn spatial hierarchies of features through multiple building blocks of convolution, pooling and fully connected layers. The convolutional layers were originally designed for image classification purposes [72], but can also be used for feature extraction from temporal sequences, including time series prediction [61]. For example, the CNN has been used for precipitation downscaling [73] and a wide time of time series prediction tasks [69]. In particular, Chandra et al. [61] showed that CNN is competitive with LSTM for multi-step ahead time series prediction. Furthermore, Hussain et al. [74] implemented a one-dimensional (1D) CNN for one-step-ahead streamflow forecasting for three-time horizons (daily, weekly and monthly) in Gilgit River of Pakistan. Khosravi et al. [75] used CNN with meta-heuristic optimisation and reported superior predictive performance in daily stream-flow forecasting compared to related models.
2.3 Ensemble learning
@@
Ensemble learning models combine multiple machine learning methods as base models to produce an optimal predictive model [39, 76]. Some of the prominent ensemble learning models are Bagging and Boosting methods via implementations such as Random Forests and AdaBoost [24], respectively. The combination of ensemble learning with deep learning models have become prominent in the last decade [39]. Ensemble models have been applied extensively in hydrology[11], with a focus on how to construct the ensemble given multiple channels of flow data using averaging [77], bootstrapping [78], and stacking [79, 80] methods. Initially, Cannon et al. [81] investigated how climate change impacted streamflow using ensemble neural networks and multiple linear regression methods. Diks an Vrugt [77] compared different model averaging techniques and indicated that the Granger-Ramanathan averaging method provided the best accuracy for streamflow forecasting. In contrast, Sharma et al. [78] demonstrated that a hybrid wavelet bootstrapped neural network provided better predictions than conventional machine learning models in forecasting daily river discharge. Tyralis et al. [79] forecasted short-term daily streamflow by stacking ten different machine learning models. Furthermore, Li et al. [80] integrated stacked models including Support Vector Machines, Elastic-net Regression and Extreme Gradient Boosting (XGBoost) to forecast midterm streamflow. However, bagging has been the predominant ensemble model until 2010 in hydrology and more recently, boosting and stacking have dominated due to improved performance [11].
2.4 Quantile Regression
Koenker et al. [30] introduced Quantile Regression for estimating point values of individual quantiles directly using regression models. Since then, quantile regression has found multiple applications in hydrology, such as precipitation downscaling [36] and climate characteristics [37]. Friederichs et al. [82] developed a censored quantile regression model to predict mixed discrete-continuous variables such as precipitation amounts, wind speeds, or pollutant concentrations. In the case of flood forecasting, Weerts et al. [38] applied quantile regression in assessing predictive uncertainty of rainfall-runoff and hydraulic forecasts.
The quantile regression model estimates the conditional median of the response variable, as compared to regular regression, which estimates the conditional mean through the least squares method. This is achieved through minimising the mean absolute error, and by applying asymmetric weights to positive/negative errors, the conditional quantiles of the response variable can be estimated instead [30]. Suppose $Y$ is the real-valued random variable for the response, then $F_{Y}(y)=P(Y\leq y)$ represents the cumulative distribution function. The $\tau^{th}$ quantile for $\tau\in(0,1)$ may be given by:
$q_{Y}(t)=F_{Y}^{-1}(\tau)=inf\{y:F_{Y}(y)\geq\tau\}$
We define the tilted absolute value function as $\rho_{\tau}(m)$ for a particular quantile $\tau$:
$\rho_{\tau}(m)=m(\tau-\mathbbm{1}_{m<0})$ (1)
where $\mathbbm{1}$ is the indicator function. The specific quantile can be found by minimising the expected loss of $Y-u$, with respect to $u$:
[…]
Quantile regression has been combined with machine learning models for several applications. Taylor et al. [33] introduced the quantile regression neural network, which offered a practical and flexible implementation compared to past work that involved linear or simple parametric non-linear models. Wang et al. [34] combined quantile regression with an LSTM model for load forecasting, and Pasche et al. [35] combined extreme value theory with quantile regression for a novel model which was applied to flood risk forecasting. These combinations motivate our work, which focuses on developing a quantile deep learning model for streamflow forecasting.